OpenStudy (anonymous):

Find equation of the line passing through (1,6) and (-1,8). Write in point slope form.

OpenStudy (nenadmatematika):

y-6=-1(x-1)

OpenStudy (anonymous):

What about slope intercept?

OpenStudy (nenadmatematika):

y=-x+7

OpenStudy (anonymous):

Standard?

OpenStudy (anonymous):

x+y=7

OpenStudy (anonymous):

Thanks!

OpenStudy (nenadmatematika):

correct :D

OpenStudy (anonymous):

m = (8-6)/ (-1-1) = 2/(-2) = -1 y -6 = -(x -1) ->y = -x +1 +6 Thus y = -x +7

OpenStudy (anonymous):

Oops, point slope form: y -6 = -(x -1)